OpenStudy (anonymous):

Hannah earned $1536 in 8 years on an investment at a 4% simple interest rate. How much was Hannah’s investment?

OpenStudy (anonymous):

@thivitaa hellp

OpenStudy (anonymous):

Didn't you just ask this? PROBLEM 1: She earns 4% simple interest each year. That's 32% after 8 years. Let x be the starting amount. The interest is computed as: 0.32x = $1536 Divide both sides by 0.32: x = 1536/0.32 x = 4800 Double-check. 4800 x 0.04 = $192 per year 8 years x $192 = $1536 Answer: Hannah started with $4800. PROBLEM 2: The price goes down by 3.50 (28 - 24.50). Compute this as a percent of the original price: 3.5 / 28 = 0.125 Turn that into a percent by moving the decimal two places to the right (multiply by 100%): Answer: The stock suffered a 12.5% decrease (or write it as -12.5%) PROBLEM 3: You made $550 interest from a $5000 investment. 550 / 5000 = 0.11 Turn it into a percent by moving the decimal two places. Answer: 11% simple interest

OpenStudy (anonymous):

I looked it up on the computer

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

You are very welcome